OpenStudy (anonymous):

convert to rectangular form r=4 cos theta

OpenStudy (nikvist):

\[r=4\cos\theta\]\[r^2=4r\cos\theta\]\[x^2+y^2=4x\]\[(x-2)^2+y^2=2^2\]

OpenStudy (anonymous):

\[ x = r \cos(\theta) \\ y = r \sin(\theta)\\ \sqrt{x^2+y^2} =r\\ r = 4 \cos(\theta)\\ \sqrt{x^2+y^2} = 4 \frac x r = 4 \frac x {\sqrt{x^2+y^2}}\\ x^2 + y^2 = 4 x \\ \]

OpenStudy (anonymous):

eliassaab, what happened to r?

OpenStudy (anonymous):

\[ r =\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

i mean he r in 4 r/x

OpenStudy (anonymous):

I have 4 x/r and not 4 r/x

OpenStudy (anonymous):

i meant that, what happened to the r

OpenStudy (anonymous):

You cross multiply to get \[ x^2 + y^2 = 4 x \]

OpenStudy (anonymous):

\[\sqrt{x^2+y^2} = 4 \frac x r = 4 \frac x {\sqrt{x^2+y^2}}\\ \sqrt{x^2+y^2} =4 \frac x {\sqrt{x^2+y^2}}\\ x^2 + y^2 = 4 x \]

OpenStudy (anonymous):

kk, does it have to be in x^2 = y^2 format to be in rectangular form?

OpenStudy (anonymous):

It has to be a relation between x and y. In our case, we have the equation of a circle.

OpenStudy (anonymous):

how would you solve r^2 = 11 cos 2 theta would you get X^2 + Y^2 = 11 x/r 2

OpenStudy (anonymous):

First you have to use the formula \[ \cos(2 \theta) = \cos^2(\theta)- sin^2(\theta) \]

OpenStudy (anonymous):

\[ x^2 + y^2 = 11 \left( \frac {x^2}{r^2}- \frac {y^2}{r^2} \right)\\ x^2 + y^2 = \frac {11}{r^2} \left( x^2- y^2 \right)\\ x^2 + y^2 = \frac {11}{x^2+y^2} \left( x^2- y^2 \right)\\ \left( x^2 + y^2 \right)^2 =11 \left( x^2- y^2 \right)\\ \]